- Performing 1550 finesse ringdown measurements first. The frequency that matches the two wavelengths has changed. The new frequency that matches them is 232.57 MHz. This is a significant drift in a 24 hour time frame.
-See attached finesse measurements. 1550ringdown.png On one trial (I took more but haven't calculated from them yet) the 1550 finesse is 2486. This is consistant with the previous cavity measurements in transmission.
On to the power budget:
-1550nm input: 12.85 mW
-1550nm REFL: 5.9 mW - This measurement is a little tricky. The has a large dependence on how well the AOM frequency is tuned. I held the power meter in the REFL spot and had brianna scan around on the aom frequency to minimize the amount read on the power meter.
-1550nm TRANS: 2.17 mW - Again, this requires placing the power meter and tuning the frequency to maximize the transmission value.
I don't think 1 - (5.9+2.17) / 12.85 = 37% worth of loss in this cavity.
-775nm input: 76 uW - note the room lights emit a fair amount of 775 light so this is done in the dark.
-775nm REFL: 75 uW make it to the REFL PD - since I can't measure the REFL power while the cavity is locked 11C98CCB-B9ED-4A2E-BF63-3207EF7A87AF.png shows the voltage value on the PD while locked. The max value is roughly 95 mV. This means nearly all the light is cancelling - (2.6 mV (dark voltage) + .8 mV )/ (2.6 mV (dark voltage) + 95 mV ) = 3% the original value. Therefore I'm inferring a power reading of 2.5 uW in reflection.
-775nm TRANS: 22 uW - after a BSW29 (50:50)
I don't think (22*2+2.5)/76 = 38% is the correct loss value.
Update Below:
I went to retry the 775 nm power measurement. I found some very high frequency oscillations in the trans signal due to the controller UGF being too high. Almost a pure sign wave. Because of this the average power the power meter is measuring is way off. This is most likely the case for the 1550 light too, will retry this tomorrow. In the mean time the new measurements are:
Input and REFL - same
TRANS - transmitted on 50:50 BS - 30.7 uW
- reflected on 50:50 BS - 36 uW
Total exiting the cavity: 66.7 uW (BSW29 on thorlabs quoted up to 1% loss or Tabs + Rabs > 99%, add 1% to this) -> 67.4 uW
This is much closer to the ringdown measurements. Tomorrow I may destroy the 775 TRANS alignment so i can measure without the BS there. And redo the 1550 measurements with this discovery in mind.
Average of 3 data points to 1550 nm ringdown in transmission. Note that there is no REFL PD to compare to (no plans to put one in either, we don't ever need to lock the cavity with 1550 light). Average finesse between the three measurements is 2580 +/- 18. Again note these error bars are only calculated using the errors in the fit calculation, other factors should be included in final numbers.